please tell me if this is correct. solve the proportion:2/11= 24/3x-4 i have 66 is this correct?
Answer by Alan3354(69443)

solve the proportion:2/11= 24/3x-4
Cross multiply
2(3x-4) = 11*24
6x-8 = 264
6x = 272
x = 136/3